Opened 6 years ago

Closed 3 months ago

#430 closed flight processing (fixed)

GB11/05, flight day 286b/2011, London

Reported by: adbe Owned by: knpa
Priority: whenever Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by adbe)

Data location: /data/okhotsk1/scratch/arsf/arsf_data/2011/flight_data/uk/GB11_05-2011_286b_London

Data arrived from ARSF via network transfer on 18/11/2011

GB11/05 Scientific objective: Atmospheric flight over London.

GB11/05 Priority: not_required

PI: Benjamin Barratt

Sensors:

Camera (23/02/2012)
LiDAR (23/02/2012)

Change History (15)

comment:1 Changed 6 years ago by adbe

Project has been split into 2, there are 2 LiDAR flights at 1500 hours, the other sensors are around 1300 hours

comment:2 Changed 6 years ago by adbe

Processing LiDAR for discrete and full waveform

comment:3 Changed 6 years ago by adbe

Processing RCD

comment:4 Changed 6 years ago by adbe

RCD ready for delivery check

comment:5 Changed 6 years ago by adbe

LiDAR ready for delivery check

comment:6 Changed 6 years ago by jaho

Checking RCD.

comment:7 Changed 6 years ago by jaho

Camera delivery check:

No issues. Ready to deliver.

comment:8 Changed 6 years ago by jaho

Checking LiDAR.

comment:9 Changed 6 years ago by jaho

LiDAR delivery check:

  • Typos in ReadMe - same as in 286b

Good otherwise. Haven't checked full waveform yet.

comment:10 Changed 6 years ago by adbe

Corrected typos in ReadMe

comment:11 Changed 6 years ago by adbe

  • Description modified (diff)

Camera and LiDAR sent to Benjamin Barratt on disk 76 on 23/02/2012

comment:12 Changed 6 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Milestone set to 2011 data processing completion
  • Owner set to knpa
  • Status changed from new to assigned

Beginning archiving.

comment:13 Changed 6 years ago by knpa

No intensity vector screenshot included with lidar delivery.

comment:14 Changed 3 months ago by asm

Archiving

All data has been archived. There was no hyperspectral data on this day. CEDA has an empty hyperspectral delivery at: http://data.ceda.ac.uk/neodc/arsf/2011/GB11_05/GB11_05-2011_286b_London/hyperspectral/

Will notify them to let them know they can delete the empty dir but no problems need resolving (or data to be uploaded).

comment:15 Changed 3 months ago by asm

  • Resolution set to fixed
  • Status changed from assigned to closed
Note: See TracTickets for help on using tickets.